OpenStudy (anonymous):

Solve t over 5 + 3 = 3.

OpenStudy (lncognlto):

So, to clarify, the question is\[\frac{ t }{ 5 + 3 } = 3\]

OpenStudy (anonymous):

t=15+9 t=24

OpenStudy (lncognlto):

So 5 + 3 = 8, so we multiply both sides of the equation by 8, like this\[\frac{ t }{ 8 } \times 8 = 3 \times \]

OpenStudy (lncognlto):

*3 x 8

OpenStudy (lncognlto):

Therefore \[t = 3 \times 8\]so\[t = ?\]